# Complexity of Abduction in Łukasiewicz Logic

### v1

@@ 1 Introduction
Abduction, deduction, and induction are three main forms of reasoning (?). Abduction (finding explanations) has multiple applications in artificial intelligence, such as diagnosis (???), commonsense reasoning (??), formalisation of scientific reasoning (?), and machine learning (?). In logic-based abduction (?), the reasoning task is to find an explanation for an observation $\chi$ from a theory $\Gamma$, i.e., a formula $\phi$ s.t. $\Gamma,\phi\models\chi$ but $\Gamma,\phi\not\models\bot$̸ ⊥ (i.e., $\Gamma$ and $\phi$ should consistently entail $\chi$).
@@
Originally (??) fuzzy logics were introduced to reason about imprecise statements such as ‘it is cold outside’, ‘the symptoms are severe’, etc. The values between $[0,1]$ are interpreted as degrees of truth from 00 (absolutely false) to $1$ (absolutely true). Fuzzy logic has also been applied to reasoning about uncertainty (cf., e.g., (?) and (?)) and beliefs (?). In knowledge representation and reasoning, fuzzy logics have found multiple applications in representing graded and fuzzy ontologies (?). In such ontologies, concept assertions and terminological axioms have degrees of truth. Moreover, fuzzy versions of description logics and their computational properties have been extensively investigated (????).
@@
The paper is structured as follows. In Section 2, we present the Łukasiewicz logic. In Section 3, we propose and motivate interval terms that we will use as solutions to Łukasiewicz abduction problems. Sections 4 and 5 are dedicated to the study of the complexity of abductive reasoning in the Łukasiewicz logic and its clausal fragment. Finally, we summarise our results and provide a plan for future work in Section 6. Omitted details of the proofs can be found in the appendix.
@@
Łϕitalic-$\mathcal{L}_{{\textsf{\L}}}\ni\phi$ $\coloneqq p\in\mathsf{Pr}\mid\neg\phi\mid(\phi\odot\phi)\mid(\phi\oplus\phi)\mid(\phi\!\rightarrow\!\phi)$
@@
$[\nicefrac{{1}}{{2}},\nicefrac{{2}}{{3}}]$ =$=\{x\mid x\in\mathbb{R},x\geq\nicefrac{{1}}{{2}},x\leq\nicefrac{{2}}{{3}}\}$
$(\nicefrac{{1}}{{2}},\nicefrac{{2}}{{3}}]_{\mathbb{Q}}$ =$=\{x\mid x\in\mathbb{Q},x>\nicefrac{{1}}{{2}},x\leq\nicefrac{{2}}{{3}}\}$
@@
We say that ϕŁitalic-$\phi\in\mathcal{L}_{{\textsf{\L}}}$ is Ł-valid (Ł${{\textsf{\L}}}\models\phi$) if $v(\phi)=1$ for every Ł-valuation $v$; $\phi$ is Ł-satisfiable if $v(\phi)=1$ for some Ł-valuation $v$.
@@
[…]
@@
Second, deduction theorem does not hold for $\models_{{\textsf{\L}}}$: indeed, while $p\models_{{\textsf{\L}}}p\odot p$, it is easy to show that Ł⊧̸${{\textsf{\L}}}\not\models p\rightarrow(p\odot p)Ł ⊧$ by setting $v(p)=\frac{1}{2}$. Third, $\odot$ and $\oplus$ are not idempotent: setting $v(p)\!=\!\frac{1}{2}$, we have that $v(p\!\odot\!p)\!=\!0$ and $v(p\!\oplus\!p)\!=\!1$.
@@
$\phi\vee\chi$ $\coloneqq(\phi\rightarrow\chi)\rightarrow\chi$ $\phi\wedge\chi$($\coloneqq\neg(\neg\phi\vee\neg\chi)$ (2)
@@ 3 Interval Terms
Before proceeding to the formal presentation of abduction, let us introduce the type of terms that we will be using in solutions. Traditionally, the form of solutions is restricted to conjunctions of literals (terms). In this case, a solution corresponds to a statement of facts. Moreover, in this case, the logically weakest solutions are the subset-minimal ones. We begin with Łukasiewicz counterparts of classical terms and clauses.
@@
$\Gamma_{\mathsf{lift}}$ =$=\left\{\begin{matrix}(c\oplus c\oplus c\oplus c)\leftrightarrow g,(\neg c\oplus\neg c\oplus\neg c)\leftrightarrow b\end{matrix}\right\}$
@@
To explain why both indicators are on, we need to present a formula $\phi$ s.t. $\Gamma_{\mathsf{lift}},\phi\models^{\mathsf{cons}}_{{\textsf{\L}}}g\odot b$. For this, we need that $v(\neg c\oplus\neg c\oplus\neg c)=1$ and $v(c\oplus c\oplus c\oplus c)=1$ in every valuation that makes $\Gamma_{\mathsf{lift}}$ true. As we noted above, this requires that $v(c)\in[\frac{1}{4},\frac{2}{3}]$. On the other hand, a simple term $\tau$ has value $1$ iff all its literals have value $1$ (i.e., all variables in $\tau$ should have value 00 or $1$). Thus, there is no simple term $\tau$ containing $c$ or $\neg c$ s.t. $\Gamma_{\mathsf{lift}},\!\tau\!\models^{\mathsf{cons}}_{{\textsf{\L}}}\!\chi$. Hence $\phi$ cannot be a simple term.
One way to circumvent this problem is to adopt the proposal of (?) and (?) and define solutions to fuzzy abduction problems as sets of assignments of values to propositional variables. This approach, however, has a drawback. In this setting, one can only express exact values of variables but not intervals of their values. Now, observe from Example 1 that any assignment of a value from $[\frac{1}{4},\frac{2}{3}]$ to $c$ solves $\langle\Gamma_{\mathsf{lift}},\chi_{\mathsf{lift}}\rangle$. Thus, in the general case, it is impossible to generate the set of all solutions as there are infinitely many of them. Moreover, it may be problematic to choose between different values.
@@
    $v(p\lozenge\mathbf{c})$ =$=\begin{cases}1&\text{if }v(p)\lozenge\mathbf{c}\\ 0&\text{otherwise}\end{cases}$ $v(\neg(p\lozenge\mathbf{c}))$ $=1-v(p\lozenge\mathbf{c})$
@@
[…]
@@
and the fact that the expansion of $\mathsf{RPL}$ with $\triangle$ is $\mathsf{NP}$-complete w.r.t. satisfiability and $\mathsf{coNP}$-complete w.r.t. validity (?), we have that the satisfiability and validity of $\mathcal{L}^{\mathbb{Q}}_{{\textsf{\L}}}$-formulas have the same complexity as those of Ł.
@@
[…]
@@ 4 Abduction in Łukasiewicz Logic
Let us now formally present abduction in Ł. We begin with the definition of abduction problems and different types of their solutions.
@@
  • An Ł-abduction problem is a tuple $\mathbb{P}=\langle\Gamma,\chi,{\mathsf{H}}\rangle$ with $\Gamma\cup\{\chi\}$ a finite set of $\mathcal{L}^{\mathbb{Q}}_{{\textsf{\L}}}$-formulas, and ${\mathsf{H}}$ a finite set of interval literals. We call $\Gamma$theory, $\chi$ an observation, and members of ${\mathsf{H}}$ hypotheses.
@@
In addition, it is easy to see that even though a theory-minimal solution is entailment-minimal, the converse is not always the case. Indeed, let =$\mathbb{P}=\langle\{p\vee q,r\},q\wedge r\rangle$. One can see that there are two entailment-minimal solutions: $p\leq\mathbf{0}$ and $q\geq\mathbf{1}$. Note, however, that $p\vee q,r,p\leq\mathbf{0}\models_{{\textsf{\L}}}q\geq\mathbf{1}$. Thus, $p\leq\mathbf{0}$ is not theory-minimal.
@@
${\mathsf{H}}_{\mathsf{lift}}$ =$=\{c\lozenge\mathbf{\tfrac{i}{12}}\mid\lozenge\in\{\leq,\geq,<,>\}\text{ and }i\in\{0,\ldots,12\}\}$
@@
$\mathcal{S}(\mathbb{P}_{\mathsf{lift}})$ =$=\left\{(c\!\triangleright\!\mathbf{\tfrac{i}{12}})\!\odot\!(c\!\triangleleft\!\mathbf{\tfrac{i^{\prime}}{12}})\left|\begin{matrix}~{}\triangleleft\!\in\!\{\leq,<\},~{}\triangleright\!\in\!\{\geq,>\},\\[3.00003pt] ~{}\mathbf{i}\!\geq\!3,~{}\mathbf{i}^{\prime}\!\leq\!8,~{}\mathbf{i}\!<\!\mathbf{i}^{\prime}\end{matrix}\right.\!\right\}\cup$
@@
[…]
@@
$p^{{\textsf{\L}}}$ $=p$ ($(\neg\phi)^{{\textsf{\L}}}$ =$=\neg\phi^{{\textsf{\L}}}$
@@
[…]
@@ 4.1 Solution Recognition
The membership follows immediately from Proposition 2 and the fact that $\tau\in\mathcal{S}(\mathbb{P})$ iff $\Gamma,\tau\models^{\mathsf{cons}}_{{\textsf{\L}}}\chi$. For the hardness, we provide a reduction from the classical solution recognition which is $\mathsf{DP}$-complete (?, §4). Let now $\mathbb{P}_{\mathsf{cl}}=\langle\Delta,\psi,{\mathsf{H}}^{\mathsf{cl}}\rangle$ be a classical abduction problem. Define Ł$\mathbb{P}^{{\textsf{\L}}}=\langle\Delta^{\sharp},\psi^{{\textsf{\L}}},{\mathsf{H}}^{\sharp}\rangle$ as follows:
@@
${\mathsf{H}}^{\sharp}$ =$=\{p\geq\mathbf{1}\mid p\in{\mathsf{H}}^{\mathsf{cl}}\}\cup\{q\leq\mathbf{0}\mid q\in{\mathsf{H}}^{\mathsf{cl}}\}$
@@
$\Gamma$ =$=\left\{p\vee\neg p\mid p\in\mathsf{Pr}(\chi\wedge q)\right\}\cup\{q\}$
${\mathsf{H}}$ =$=\{p\geq\mathbf{1}\mid p\in\mathsf{Pr}(\chi)\}\cup\{p\leq\mathbf{0}\mid p\in\mathsf{Pr}(\chi)\}$
@@
Now, we use an $\mathsf{NP}$-oracle that guesses linearly many w.r.t. $|{\mathsf{H}}|$-valuations $v_{\mathsf{sat}}$, $v_{\mathsf{prp}}$, and $v_{\lambda}$ (for each $\lambda\in\tau$) and verifies whether they witness (i) $\Gamma,\tau\not\models_{{\textsf{\L}}}\bot$, (ii) $\tau\not\models_{{\textsf{\L}}}\chi$, and (iii) $\Gamma,\tau^{\flat}_{\lambda}\not\models_{{\textsf{\L}}}\chi$. Parallel to that (as we do not need the results of (i)–(iii)), we conduct a $\mathsf{coNP}$ check that (iv) $\Gamma,\tau\models_{{\textsf{\L}}}\chi$. It follows from the definition of $\models_{{\textsf{\L}}}$-minimal solutions that $\tau$ is a $\models_{{\textsf{\L}}}$-minimal solution iff the $\mathsf{NP}$ and $\mathsf{coNP}$ checks succeed. ∎
@@ 4.2 Solution Existence
$\Gamma_{\mathsf{cl}}$ =$=\{\!\neg\phi\!\vee\!(p\!\wedge\!\tau),\!\neg p\!\vee\!\tau\}\!\cup\!\{\neg r\!\Leftrightarrow\!r^{\prime}\!\mid\!r\!\in\!\mathsf{Pr}(\phi)\!\setminus\!\mathsf{Pr}(p\!\wedge\!\tau)\!\}$ ($p\notin\mathsf{Pr}(\phi\wedge\tau)$, $\tau$ is a weak conjunction of literals)
@@
${\mathsf{H}}$ =$=\{r\mid r\!\in\!\mathsf{Pr}(\phi)\!\setminus\!\mathsf{Pr}(p\!\wedge\!\tau)\}\!\cup\!\{r^{\prime}\mid\neg r\!\Leftrightarrow\!r^{\prime}\!\in\!\Gamma_{\mathsf{cl}}\}$ (4)
By (?, Theorem 4.2), determining the existence of classical solutions for these problems is $\Sigma^{\mathsf{P}}_{2}$-hard. We reduce $\mathbb{P}_{\mathsf{cl}}$ to Ł$\mathbb{P}^{{\textsf{\L}}}=\langle\Gamma^{\sharp},\chi^{{\textsf{\L}}}_{\mathsf{cl}},{\mathsf{H}}^{\sharp}\rangle$ as was done in Theorem 1 (recall Definition 6 for $\Gamma^{{\textsf{\L}}}$ and $\chi^{{\textsf{\L}}}$) with
@@
${\mathsf{H}}^{\sharp}$ =$=\{s\geq\mathbf{1}\mid s\in{\mathsf{H}}\}\quad\chi^{{\textsf{\L}}}_{\mathsf{cl}}=p\odot\bigodot\limits_{l\in\tau}l$ (5)

### v3

@@ 1 Introduction
Abduction, deduction, and induction are three main forms of reasoning (?). Abduction (finding explanations) has multiple applications in artificial intelligence, such as diagnosis (???), commonsense reasoning (??), formalisation of scientific reasoning (?), and machine learning (?). In logic-based abduction (?), the reasoning task is to find an explanation for an observation $\chi$ from a theory $\Gamma$, i.e., a formula $\phi$ s.t. $\Gamma,\phi\models\chi$ but $\Gamma,\phi\not\models\bot$ (i.e., $\Gamma$ and $\phi$ should consistently entail $\chi$).
@@
Originally (??) fuzzy logics were introduced to reason about imprecise statements such as ‘it is cold outside’, ‘the symptoms are severe’, etc. The values between $[0,1]$ are interpreted as degrees of truth from 0 (absolutely false) to $1$ (absolutely true). Fuzzy logic has also been applied to reasoning about uncertainty (cf., e.g., (?) and (?)) and beliefs (?). In knowledge representation and reasoning, fuzzy logics have found multiple applications in representing graded and fuzzy ontologies (?). In such ontologies, concept assertions and terminological axioms have degrees of truth. Moreover, fuzzy versions of description logics and their computational properties have been extensively investigated (????).
@@
The paper is structured as follows. In Section 2, we present the Łukasiewicz logic. In Section 3, we propose and motivate interval terms that we will use as solutions to Łukasiewicz abduction problems. Sections 4 and 5 are dedicated to the study of the complexity of abductive reasoning in the Łukasiewicz logic and its clausal fragment. Finally, we summarise our results and provide a plan for future work in Section 6. Due to limited space, some proofs are put into the appendix of the full version (?).
@@
$\mathcal{L}_{{\textsf{\L }}}\ni\phi$ $\coloneqq p\in\mathsf{Pr}\mid\neg\phi\mid(\phi\odot\phi)\mid(\phi\oplus\phi)\mid(\phi\!\rightarrow\!\phi)$
@@
$[\nicefrac{{1}}{{2}},\nicefrac{{2}}{{3}}]$ $=\{x\mid x\in\mathbb{R},x\geq\nicefrac{{1}}{{2}},x\leq\nicefrac{{2}}{{3}}\}$
$(\nicefrac{{1}}{{2}},\nicefrac{{2}}{{3}}]_{\mathbb{Q}}$ $=\{x\mid x\in\mathbb{Q},x>\nicefrac{{1}}{{2}},x\leq\nicefrac{{2}}{{3}}\}$
@@
We say that $\phi\in\mathcal{L}_{{\textsf{\L }}}$ is Ł-valid (${{\textsf{\L }}}\models\phi$) if $v(\phi)=1$ for every Ł-valuation $v$; $\phi$ is Ł-satisfiable if $v(\phi)=1$ for some Ł-valuation $v$.
@@
$\phi\equiv_{{\textsf{\L }}}\chi$ $\text{ iff }\forall v:v(\phi)=v(\chi)$
$\phi\simeq_{{\textsf{\L }}}\chi$ $\text{ iff }\forall v:v(\phi)=1\leftrightharpoons v(\chi)=1$
@@
Second, deduction theorem does not hold for $\models_{{\textsf{\L }}}$: indeed, while $p\models_{{\textsf{\L }}}p\odot p$, it is easy to show that ${{\textsf{\L }}}\not\models p\rightarrow(p\odot p)$ by setting $v(p)=\frac{1}{2}$. Third, $\odot$ and $\oplus$ are not idempotent: setting $v(p)\!=\!\frac{1}{2}$, we have that $v(p\!\odot\!p)\!=\!0$ and $v(p\!\oplus\!p)\!=\!1$.
@@
$\phi\vee\chi$ $\coloneqq(\phi\rightarrow\chi)\rightarrow\chi$ $\phi\wedge\chi$ $\coloneqq\neg(\neg\phi\vee\neg\chi)$ (2)
@@ 3 Interval Terms
Before proceeding to the formal presentation of abduction, let us introduce the terms that we will be using in solutions. Traditionally, the form of solutions is restricted to conjunctions of literals (terms). In this case, a solution corresponds to a statement of facts. Moreover, in this case, the logically weakest solutions are the subset-minimal ones. We begin with Łukasiewicz counterparts of classical terms and clauses.
@@
$\Gamma_{\mathsf{lift}}$ $=\left\{\begin{matrix}(c\oplus c\oplus c\oplus c)\leftrightarrow g,(\neg c\oplus\neg c\oplus\neg c)\leftrightarrow b\end{matrix}\right\}$
@@
To explain why both indicators are on, we need to present a formula $\phi$ s.t. $\Gamma_{\mathsf{lift}},\phi\models^{\mathsf{cons}}_{{\textsf{\L }}}g\odot b$. For this, we need that $v(\neg c\oplus\neg c\oplus\neg c)=1$ and $v(c\oplus c\oplus c\oplus c)=1$ in every valuation that makes $\Gamma_{\mathsf{lift}}$ true. As we noted above, this requires that $v(c)\in[\frac{1}{4},\frac{2}{3}]$. On the other hand, a simple term $\tau$ has value $1$ iff all its literals have value $1$ (i.e., all variables in $\tau$ should have value 0 or $1$). Thus, there is no simple term $\tau$ containing $c$ or $\neg c$ s.t. $\Gamma_{\mathsf{lift}},\!\tau\!\models^{\mathsf{cons}}_{{\textsf{\L }}}\!\chi$. Hence $\phi$ cannot be a simple term.
One way to circumvent this problem is to adopt the proposal of (?) and (?) and define solutions to fuzzy abduction problems as sets of assignments of values to propositional variables. This approach, however, has a drawback. In this setting, one can only express exact values of variables but not intervals of their values. Now, observe from Example 1 that any assignment of a value from $\frac{1}{4},\frac{2}{3}]$ to $c$ solves $\langle\Gamma_{\mathsf{lift}},\chi_{\mathsf{lift}}\rangle$. Thus, in the general case, it is impossible to generate the set of all solutions as there are infinitely many of them. Moreover, it may be problematic to choose between different values.
@@
    $v(p\lozenge\mathbf{c})$ $=\begin{cases}1&\text{if }v(p)\lozenge\mathbf{c}\\ 0&\text{otherwise}\end{cases}$ $v(\neg(p\lozenge\mathbf{c}))$ $=1-v(p\lozenge\mathbf{c})$
@@
$p\leq\mathbf{c}\models_{{\textsf{\L }}}p\leq\mathbf{c}^{\prime}$ iff $𝐜\displaystyle\text${ iff $\mathbf$$\leq\mathbf{c}^{\prime}$ $p\geq\mathbf{c}\models_{{\textsf{\L }}}p\geq\mathbf{c}^{\prime}$ iff $𝐜\displaystyle\text${ iff $\mathbf$$\geq\mathbf{c}^{\prime}$
@@
and expanding the constraint tableaux calculus of (?) with rules for $\triangle$ and rational constants, we have that the satisfiability and validity of $\mathcal{L}^{\mathbb{Q}}_{{\textsf{\L }}}$-formulas have the same complexity as those of Ł.
@@
We note briefly that interval terms allow us to express not only the values of variables but also of arbitrary formulas (cf. (?) for an alternative approach). For $\phi\!\in\!\mathcal{L}_{{\textsf{\L }}}$ and $p\!\in\!\mathsf{Pr}$ s.t. $p\!\notin\!\mathsf{Pr}(\phi)$, $c\!\in\![0,1]_{\mathbb{Q}}$, and ${,<,,>}\lozenge\!\in\!\{\leq$$\geq,$>\}, we have $v(\phi)\lozenge c$ iff $v((\phi\leftrightarrow p)\odot p\lozenge\mathbf{c})=1$.
@@ 4 Abduction in Łukasiewicz Logic
Let us now present abduction in Ł. Our idea is to use interval terms as solutions to problems $\mathbb{P}=\langle\Gamma,\chi,{\mathsf{H}}\rangle$. Here, $\mathsf{H}$} is the set of hypotheses (interval literals) that one can use to build solutions. One can restrict it in two ways. First, one may allow arbitrary interval literals over a given finite set of variables. Second, one can explicitly define a finite set of interval literals. We choose the second option, as the first one leads to infinite sets of solutions (cf. Example 1).
@@
[…]
@@
In addition, it is easy to see that even though a theory-minimal solution is entailment-minimal, the converse is not always the case. Indeed, let $\mathbb{P}=\langle\{p\vee q,r\},q\wedge r\rangle$. One can see that there are two entailment-minimal solutions: $p\leq\mathbf{0}$ and $q\geq\mathbf{1}$. Note, however, that $p\vee q,r,p\leq\mathbf{0}\models_{{\textsf{\L }}}q\geq\mathbf{1}$. Thus, $p\leq\mathbf{0}$ is not theory-minimal.
@@
${\mathsf{H}}_{\mathsf{lift}}$ $=\{c\lozenge\mathbf{\tfrac{i}{12}}\mid\lozenge\in\{\leq,\geq,<,>\}\text{ and }i\in\{0,\ldots,12\}\}$
@@
[…]
@@
Let $\mathcal{L}_{\mathsf{CPL}}$ be the propositional language over $\{\neg,\wedge,\vee\}$. A classical abduction problem is a tuple $\mathbb{P}=\langle\Gamma,\chi,{\mathsf{H}}\rangle$ s.t. $\Gamma\!\cup\!\{\chi\}\!\subseteq\!\mathcal{L}_{\mathsf{CPL}}$ and $\mathsf{H}$} is a set of simple literals.
@@
$p^{{\textsf{\L }}}$ $=p$ $(\neg\phi)^{{\textsf{\L }}}$ $=\neg\phi^{{\textsf{\L }}}$
@@
$\Gamma\models_{\mathsf{CPL}}\chi$ $\text{ iff }\Gamma^{{\textsf{\L }}},\{p\vee\neg p\mid p\in\mathsf{Pr}[\Gamma\cup\{\chi\}]\}\models_{{\textsf{\L }}}\chi^{{\textsf{\L }}}$
@@ 4.1 Solution Recognition
The membership follows immediately from Proposition 2 and the fact that $\tau\in\mathcal{S}(\mathbb{P})$ iff $\Gamma,\tau\models^{\mathsf{cons}}_{{\textsf{\L }}}\chi$. For the hardness, we provide a reduction from the classical solution recognition which is $\mathsf{DP}$-complete (?, §4). Let now $\mathbb{P}_{\mathsf{cl}}=\langle\Delta,\psi,{\mathsf{H}}^{\mathsf{cl}}\rangle$ be a classical abduction problem. Define $\mathbb{P}^{{\textsf{\L }}}=\langle\Delta^{\sharp},\psi^{{\textsf{\L }}},{\mathsf{H}}^{\sharp}\rangle$ as follows:
@@
${\mathsf{H}}^{\sharp}$ $=\{p\geq\mathbf{1}\mid p\in{\mathsf{H}}^{\mathsf{cl}}\}\cup\{q\leq\mathbf{0}\mid q\in{\mathsf{H}}^{\mathsf{cl}}\}$
@@
$\Gamma$ $=\left\{p\vee\neg p\mid p\in\mathsf{Pr}(\chi\wedge q)\right\}\cup\{q\}$
${\mathsf{H}}$ $=\{p\geq\mathbf{1}\mid p\in\mathsf{Pr}(\chi)\}\cup\{p\leq\mathbf{0}\mid p\in\mathsf{Pr}(\chi)\}$
@@
Now, we use an $\mathsf{NP}$-oracle that guesses linearly many w.r.t. $|{\mathsf{H}}|$ Ł-valuations $v_{\mathsf{sat}}$, $v_{\mathsf{prp}}$, and $v_{\lambda}$ (for each $\lambda\in\tau$) and verifies whether they witness (i) $\Gamma,\tau\not\models_{{\textsf{\L }}}\bot$, (ii) $\tau\not\models_{{\textsf{\L }}}\chi$, and (iii) $\Gamma,\tau^{\flat}_{\lambda}\not\models_{{\textsf{\L }}}\chi$. Parallel to that (as we do not need the results of (i)–(iii)), we conduct a $\mathsf{coNP}$ check that (iv) $\Gamma,\tau\models_{{\textsf{\L }}}\chi$. It follows from the definition of $\models_{{\textsf{\L }}}$-minimal solutions that $\tau$ is a $\models_{{\textsf{\L }}}$-minimal solution iff the $\mathsf{NP}$ and $\mathsf{coNP}$ checks succeed. ∎
@@ 4.2 Solution Existence
$\Gamma_{\mathsf{cl}}$ $=\{\!\neg\phi\!\vee\!(p\!\wedge\!\tau),\!\neg p\!\vee\!\tau\}\!\cup\!\{\neg r\!\Leftrightarrow\!r^{\prime}\!\mid\!r\!\in\!\mathsf{Pr}(\phi)\!\setminus\!\mathsf{Pr}(p\!\wedge\!\tau)\!\}$ ($p\notin\mathsf{Pr}(\phi\wedge\tau)$, $\tau$ is a weak conjunction of literals)
@@
${\mathsf{H}}$ $=\{r\mid r\!\in\!\mathsf{Pr}(\phi)\!\setminus\!\mathsf{Pr}(p\!\wedge\!\tau)\}\!\cup\!\{r^{\prime}\mid\neg r\!\Leftrightarrow\!r^{\prime}\!\in\!\Gamma_{\mathsf{cl}}\}$ (4)
By (?, Theorem 4.2), determining the existence of classical solutions for these problems is $\Sigma^{\mathsf{P}}_{2}$-hard. We reduce $\mathbb{P}_{\mathsf{cl}}$ to $\mathbb{P}^{{\textsf{\L }}}=\langle\Gamma^{\sharp},\chi^{{\textsf{\L }}}_{\mathsf{cl}},{\mathsf{H}}^{\sharp}\rangle$ as was done in Theorem 1 (recall Definition 6 for $\Gamma^{{\textsf{\L }}}$ and $\chi^{{\textsf{\L }}}$) with
@@
${\mathsf{H}}^{\sharp}$ $=\{s\geq\mathbf{1}\mid s\in{\mathsf{H}}\}\quad\chi^{{\textsf{\L }}}_{\mathsf{cl}}=p\odot\bigodot\limits_{l\in\tau}l$ (5)